Cochrane, Inc., is considering a new three-year expansion project that requires an initial fixed asset investment of $2.1 million. The fixed asset will be depreciated straight-line to zero over its three-year tax life, after which time it will be worthless. The project is estimated to generate $2,150,000 in annual sales, with costs of $1,078,327. Required: If the tax rate is 35 percent, what is the OCF for this project? (Do not include the dollar sign ($). Enter your answer in dollars(e.g., 1,234,567), not millions of dollars.).
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(2,100,000/3)=$700,000/year

OCF=(Sales-Costs)(1-tax rate)+Tax savings on Annual depreciation

=(2,150,000-1,078,327)(1-0.35)+(0.35*700,000)

=941587.45